Grade 12th passMechanics

At what angle a pratical is projected so that its height attained is equal to its horizontal range

Biplob Mondol
8 Years agoGrade 12th pass

1 Answer

NIKHIL GHODKE
8 Years ago
 we know the formula for max height and range height =(u sin(o))2/2grange=u2sin(2o) /g
also
sin(2o)=2sin(o)cos(o)
from rang eis equal to height we get
sin(o)/2=sin(2o)
sin(o)/2=2sin(o)cos(o)
sin(o)=4cos(o)
tan(o)=4